Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by Anindya Madhudor » Thu Dec 13, 2012 8:49 am
A box contains three pairs of blue gloves and two pairs of green gloves. Each pair consists of a left-hand glove and a right-hand glove. Each of the gloves is separate from its mate and thoroughly mixed together with the others in the box. If three gloves are randomly selected from the box, what is the probability that a matched set (i.e., a left- and right-hand glove of the same color) will be among the three gloves selected?

a. 3/10
b. 23/60
c. 7/12
d. 41/60
e. 5/6

OA: D
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Thu Dec 13, 2012 6:42 pm
Anindya Madhudor wrote:A box contains three pairs of blue gloves and two pairs of green gloves. Each pair consists of a left-hand glove and a right-hand glove. Each of the gloves is separate from its mate and thoroughly mixed together with the others in the box. If three gloves are randomly selected from the box, what is the probability that a matched set (i.e., a left- and right-hand glove of the same color) will be among the three gloves selected?

a. 3/10
b. 23/60
c. 7/12
d. 41/60
e. 5/6

OA: D
P(matching set) = 1 - P(no matching set).

There are 3 ways to NOT select a matching set of gloves.

Case 1: 3 non-matching blue
P(1st glove is blue) = 6/10.
P(2nd glove is blue and for the same hand) = 2/9.
P(3rd glove is blue and for the same hand) = 1/8.
Since we want all of these events to happen, we multiply the fractions:
6/10 * 2/9 * 1/8 = 1/60.

Case 2: 1 green, 2 non-matching blue
P(1st glove is green) = 4/10.
P(2nd glove is blue) = 6/9.
P(3rd glove is blue and for the same hand as the 2nd glove) = 2/8.
Since we want all of these events to happen, we multiply the fractions:
4/10 * 6/9 * 2/8 = 4/60.
Since the one green glove could be 1st, 2nd or 3rd, we multiply by 3:
4/60 * 3 = 12/60.

Case 3: 1 blue, 2 non-matching green
P(1st glove is blue) = 6/10.
P(2nd glove is green) = 4/9.
P(3rd glove is green and for the same hand as the 2nd glove) = 1/8.
Since we want all of these events to happen, we multiply the fractions:
6/10 * 4/9 * 1/8 = 1/30.
Since the one blue glove could be 1st, 2nd or 3rd, we multiply by 3:
1/30 * 3 = 3/30 = 6/60.

Since any of the 3 cases above would yield no matching set of gloves, we add the probabilities:
P(no matching set) = 1/60 + 12/60 + 6/60 = 19/60.

Thus, P(matching set) = 1 - 19/60 = 41/60.

The correct answer is D.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Anindya Madhudor » Fri Dec 14, 2012 5:18 am
Thank you, Mitch. You are awesome.
Join the discussion

by viveksingh222 » Fri Dec 14, 2012 10:55 pm
Why this question can't be answered by below method..
Total number of possible combinations = 10C3 = 120.
No of ways to pick 3 such that a pair is included is 5C1*8C1(5C1= selection of a pair from 5, and 8C1= selection of 1 glove from remaining 8)

=80

Thus the probability of selecting a pair would be 80/120 or 40/60..

but still I am missing the exact answer 41/60...
can you please explain where i am actually missing or this isn't the right approach to be followed for solving this question..
Thanks....
Join the discussion

by GMATGuruNY » Sat Dec 15, 2012 4:59 am
viveksingh222 wrote:Why this question can't be answered by below method..
Total number of possible combinations = 10C3 = 120.
No of ways to pick 3 such that a pair is included is 5C1*8C1(5C1= selection of a pair from 5, and 8C1= selection of 1 glove from remaining 8)

=80

Thus the probability of selecting a pair would be 80/120 or 40/60..

but still I am missing the exact answer 41/60...
can you please explain where i am actually missing or this isn't the right approach to be followed for solving this question..
Thanks....
The portion in red assumes that there are 5 DISTINCT pairs of gloves.
Let LB = left-handed blue glove, RB = right-handed blue glove, LG = left-handed green glove, and RG = right-handed green glove.
The 3 pairs of blue gloves and 2 pairs of green gloves are as follows:
LB, LB, LB, RB, RB, RB, LG, LG, RG, RG.
Any LB can be combined with any RB to form a matched set of blue gloves.
Either LG can be combined with either RG to form a matched set of green gloves.
Thus, there are more than 5 ways to choose a matched set of gloves.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by viveksingh222 » Sat Dec 15, 2012 5:19 am
hmm..it makes sense now...thanks...
Join the discussion

by ritind » Mon Dec 17, 2012 2:12 am
Mitch, i'm trying to use combination technique bt its nt working here. can you explain thru dis method
Join the discussion